# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 4 Joint Tracking Error and Weight Constraints
with
$\hat{\boldsymbol{l}}=\frac{\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}}}{\hat{w}_{k}-\hat{w}_{a}},$
and
$\hat{\boldsymbol{k}}=\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}},\quad\hat{\boldsymbol{a}}=\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}},$
with
$\hat{w}_{k}-\hat{w}_{a}=\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}},$
and $\hat{w_{k}}\neq\hat{w}_{a}$ since by assumption all assets in the portfolio cannot be restricted. We also have
$\hat{w}_{u}=\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{\mu}}-\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}.$
First define the restricted benchmark portfolio as $\boldsymbol{m}_{R}$ as the benchmark portfolio. Define $\boldsymbol{w}_{R}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$ and
$\hat{\boldsymbol{w}}_{R}\mathrel{\mathop{:}}=\hat{\boldsymbol{w}}_{cp}+\boldsymbol{m}_{R}.$ (10)
Define now the Sharpe Ratio associated with joint restrictions of TE and weight:
$SR_{R}\mathrel{\mathop{:}}=\frac{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu}}{\sqrt{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}},$
and estimate
$\widehat{SR}_{R}\mathrel{\mathop{:}}=\frac{\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\mu}}{\sqrt{\hat{\boldsymbol{w}}_{R}}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}.$
We need the following sparsity assumption which is strenghtened version in terms of sparsity compared with Assumption 7.
Assumption 8.
(i). $\bar{s}^{6}l_{T}r_{T}^{5}K^{23/2}=o(1)$.
(ii). $|\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{p}|\geq c>0,|\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}|\geq c>0$, $|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}|\geq c>0$, with $c$ being a positive constant.
(iii). $10<r/p<$ 1, $|w_{k}-w_{a}|\geq c>0$, with $c$ being a positive constant, and $|\omega|\leq C<\infty,$ $0<c\leq|\kappa|\leq C<\infty.$
@@
Assumption 8(i) replaces Assumption 7(i) and shows the difficulty of joint TE and weight constraints. Assumption 8(i) rate is slower compared with Assumption 7(i), but can be satisfied with the following example. Let $\bar{s}=ln(T),K=ln(T),r_{T}=T^{1/20},p=2T,\lambda=\sqrt{ln(2T)/T}$ and $l_{T}$ definition in (2) is
$l_{T}=T^{2/20}(lnT)^{5/2}(lnT)\sqrt{ln(2T)/T}=O(\frac{1}{T^{8/20}}(lnT)^{4}).$
Then Assumption 8(i) rate is
$\bar{s}^{6}r_{T}^{5}l_{T}K^{23/2}=O(\frac{(lnT)^{6}(lnT)^{4}T^{5/20}}{T^{8/20}}(lnT)^{23/2})=o(1).$
Assumption 8(ii) provides a lower bound for returns scaled by variance, and they cannot be zero and cannot converge to zero and also the optimal portfolio variance has to be positive, and the restricted version variance has to be positive, cannot be zero or converge to zero, and this is mainly discussed also after Assumption 7. Assumption 8(iii) assumes the number of restrictions can be at least 1, and cannot be as large as $p$, so we cannot restrict all the portfolio, but we can restrict $p-1$ assets at most. Technical term $w_{k}\neq w_{a}$ since we cannot have full restrictions $p$, but here we assume that its absolute value cannot converge to zero as well. Our equality restrictions cannot be unbounded and has to be bounded by a positive finite constant in absolute value, and our risk tolerance parameter in absolute terms cannot be unbounded, and cannot converge to zero either, and has to be upper bounded by a positive constant. Assumption 8(iv) also does not allow restricted portfolio return to be zero (nor converge to zero), and the variance cannot converge to zero, and the constrained benchmark portfolio, has to have the same order as the optimal joint TE and weight portfolio, $w_{cp}^{*}$ in the optimization at the start of this section.
We have the following theorem that shows it is possible to estimate a portfolio with weight and tracking error constraints in large dimensions. Also consistent estimation of Sharpe ratio is possible and we allow for $p>n$.
Theorem 2.
@@
(i). Define $r_{w2}\mathrel{\mathop{:}}=O(\bar{s}^{4}r_{T}^{3}l_{T}K^{11/2})$, then
$\|\hat{\boldsymbol{w}}_{R}-\boldsymbol{w}_{R}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*}\|_{1}=O_{p}(r_{w2})=o_{p}(1).$
@@
$\|\frac{\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}-1\|_{1}=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6})=o_{p}(1).$
@@
$\left|\left(\frac{\widehat{SR}_{R}}{SR_{R}}\right)^{2}-1\right|=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6})=o_{p}(1).$
Remarks. 1.Comparing Theorem 1 with Theorem 2 shows that the rates with joint tracking and weight constraints are slower. This shows the negative impact of the weight constraints when added with tracking error constraints.
@@
2. We can also have an estimate for tracking error as $\widehat{TE}_{cp}\mathrel{\mathop{:}}=\sqrt{\hat{\boldsymbol{w}}_{cp}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{cp}}$, as an out-sample estimate. It can be shown that $|(\frac{\widehat{TE}_{cp}}{TE})^{2}-1|$ will converge in probability to zero by following the proof of Theorem 2(ii), and the rate will be the same as in Theorem 2(ii). This estimator is consistent.
@@
3. There is also an issue of how TE constraints jointly with weight constraints affect the returns. Information Ratio (IR) is used to measure the effect of TE constraints on returns, however Bajeux-Besnainou et al. (2011) shows that IR in case of weights joint with TE constraints is not an appropriate measure. The reason IR is not appropriate since it can decrease artificially with an increase in TE. They propose Adjusted Information Ratio (AIR). In (22) of Bajeux-Besnainou et al. (2011), it is clear that AIR is measuring the returns of optimal portfolio minus constrained minimum TE portfolio satisfying the weight constraints with respect to risk of that portfolio. We will not be pursuing such approach, since our interest centers on optimal portfolios as in Theorem 2.
4. The case of only weight constraints are analyzed after the proof of Theorem 2 in Appendix, and the rates are the same as in Theorem 2.
5 Tracking Error with Inequality Constraints on the Weights of the Portfolio
@@
In practice, inequality constraints on the weight of the portfolio is used in practice a lot. Of course, one major issue is that how the inequality weight interacts with large number of assets in the portfolio? In this part, we assume $\omega\geq 0$, by that restriction we exclude the case where benchmark portfolio, $m$, does not satisfy the inequality constraint. So we exclude an unlikely event in practice.
Optimization is:
@@
Note that when the weight constraint is not binding, we should get Tracking Error (TE) solution $\boldsymbol{w}_{d}^{*}$ in Section 2. Otherwise, we should get the solution in Section 3, which is $\boldsymbol{w}_{cp}^{*}$. This $\boldsymbol{w}_{cp}^{*}$ is the joint tracking error and equality weight constraint based portfolio. We analyze the cases that may lead to binding or nonbinding weight constraint solution now. To have a binding solution $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}^{*}\geq\omega$ which is from section 3. Or we can represent this solution as by $\boldsymbol{w}_{d}^{*},w_{u}$ definitions $\kappa w_{u}\geq\omega$ since $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}^{*}=\kappa w_{u}$ by (3)(8).  [^1] As in Bajeux-Besnainou et al. (2011) we use greater than equal constraint for binding rather than just equality in section 3. The nonbinding case will be $\kappa w_{u}<\omega$, as also seen in (17) of Bajeux-Besnainou et al. (2011).
We provide the following theoretical portfolio
$\boldsymbol{w}_{op}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}_{d}^{*}1_{\{\kappa w_{u}<\omega\}}+\boldsymbol{w}_{cp}^{*}1_{\{\kappa w_{u}>\omega\}}.$
@@
Basically, this is the portfolio if the weight constraint is not binding, i.e. $\kappa w_{u}<\omega$, then the portfolio is $\boldsymbol{w}_{d}^{*}$ which is TE constrained portfolio, and if the weight constraint is binding $\kappa w_{u}>\omega$, then the portfolio is given by joint TE and weight constraint portfolio which is $\boldsymbol{w}_{cp}^{*}$. For technical reasons we exclude $\kappa w_{u}=\omega$.The feasible estimator of this theoretical portfolio is given by
$\hat{\boldsymbol{w}}_{op}\mathrel{\mathop{:}}=\hat{\boldsymbol{w}}_{d}1_{\{\kappa\hat{w}_{u}<\omega\}}+\hat{\boldsymbol{w}}_{cp}1_{\{\kappa\hat{\boldsymbol{w}}_{u}>\omega\}}.$
Note that the estimated portfolio with benchmarks taken into account is defined as by (4)(10)
$\hat{\boldsymbol{w}}_{est}\mathrel{\mathop{:}}=\hat{\boldsymbol{w}}1_{\{\kappa\hat{\boldsymbol{w}}_{u}<\omega\}}+\hat{\boldsymbol{w}}_{R}1_{\{\kappa\hat{\boldsymbol{w}}_{u}>\omega\}}.$
@@
The theoretical counterpart is given by 222 An equivalent way of writing the nonbinding constraint in the indicator function is by $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}<\omega+\boldsymbol{1}_{R}^{\prime}\boldsymbol{m}$, and the binding constraint in the indicator function can be written as $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}>\omega+\boldsymbol{1}_{R}^{\prime}\boldsymbol{m}$
$\boldsymbol{w}_{est}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}^{*}1_{\{\kappa w_{u}<\omega\}}+\boldsymbol{w}_{R}^{*}1_{\{\kappa w_{u}>\omega\}}.$
We define the Sharpe Ratio whether we are in the binding weights case or nonbinding case as follows:
$SR_{est}^{*}\mathrel{\mathop{:}}=\frac{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu}}{\sqrt{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}}1_{\{\kappa w_{u}<\omega\}}+\frac{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu}}{\sqrt{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}}1_{\{\kappa w_{u}>\omega\}},$
and its estimator as
$\widehat{SR}_{est}\mathrel{\mathop{:}}=\frac{\hat{\boldsymbol{w}}^{\prime}\mu}{\sqrt{\hat{\boldsymbol{w}}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}}1_{\{\kappa\hat{w}_{u}<\omega\}}+\frac{\hat{\boldsymbol{w}}_{R}^{{}^{\prime}}\boldsymbol{\mu}}{\sqrt{\hat{\boldsymbol{w}}_{R}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}}1_{\{\kappa\hat{w}_{u}>\omega\}}.$
@@
Next theorem shows our weight, variance and Sharpe Ratio estimates are consistent, regardless of inequality constraints are binding or not. This is a new result in the literature. In reality, we do not know whether the constraints are binding or not, but our data based method selects whether the constraints are binding or not and our estimates converge in probability to the optimal weight, variance or Sharpe-Ratio. Let $c>0$ be a positive constant.
Theorem 3.
@@
(i). If $\kappa w_{u}<\omega$, (non-binding weight restriction) then
$\|\hat{\boldsymbol{w}}_{est}-\boldsymbol{w}_{est}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{op}-\boldsymbol{w}_{op}^{*}\|_{1}=O_{p}(r_{w1})=o_{p}(1),$
but if $\kappa w_{u}>\omega$ (binding weight restriction)
$\|\hat{\boldsymbol{w}}_{est}-\boldsymbol{w}_{est}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{op}-\boldsymbol{w}_{op}^{*}\|_{1}=O_{p}(r_{w2})=o_{p}(1).$

### v2

@@ 4 Joint TE and Weight Constraints
with $\hat{\boldsymbol{l}}=\frac{\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}}}{\hat{w}_{k}-\hat{w}_{a}}$,   and $\hat{\boldsymbol{k}}=\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}},\quad\hat{\boldsymbol{a}}=\frac{\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}$, with $\hat{w}_{k}-\hat{w}_{a}=\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}},$ and $\hat{w_{k}}\neq\hat{w}_{a}$ since, by assumption, not all assets in the portfolio are restricted. We also have $\hat{w}_{u}=\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{\mu}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{\mu}}-\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}.$ To distinguish from the general benchmark $\boldsymbol{m}$ we denote $\boldsymbol{m}_{R}$ as the restricted benchmark portfolio. Define $\boldsymbol{w}_{R}^{*}:=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$, and
$\hat{\boldsymbol{w}}_{R}:=\hat{\boldsymbol{w}}_{cp}+\boldsymbol{m}_{R}.$ (10)
Define the Sharpe Ratio associated with joint restrictions of TE and weight: $SR_{R}:=\frac{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu}}{\sqrt{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}},$ and estimate $\widehat{SR}_{R}:=\frac{\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\mu}}{\sqrt{\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}}.$ We need the following assumption which puts a lower bound on restricted portfolio terms.
Assumption 8
(i). $\bar{s}l_{T}K^{4}=o(1)$.
(ii). $|\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{p}|\geq c>0,\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}\geq c>0$, $|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}|\geq c>0$, with $c,C$ being positive constants.
(iii). $0<r/p<1$, $|w_{k}-w_{a}|\geq c>0$, with $c$ being a positive constant, and $|\omega|\leq C<\infty,$ $0<c\leq|\kappa|\leq C<\infty.$
@@
Assumption 8(i) is the same as Assumption 7(i). Assumption 8(ii) provides a lower bound for returns scaled by variance, and they cannot be zero or converge to zero. Global Minimum Variance Portfolio variance has to be finite. The restricted version variance has to be finite. Assumption 8(iii) assumes the number of restrictions can be at least 1, and cannot be as large as $p$, so we can restrict p1p-1 assets at most. Technical term $w_{k}\neq w_{a}$ since we cannot have full restrictions on $p$ assets, but here we assume that $|w_{k}-w_{a}|$ cannot converge to zero as well. Our equality restrictions have to be bounded by a positive finite constant in absolute value, and our risk tolerance parameter in absolute terms cannot be unbounded. It cannot converge to zero either, and has to be upper bounded by a positive constant. Assumption 8(iv) also does not allow restricted portfolio return to be zero (nor converge to zero), the variance cannot converge to zero, and the constrained benchmark portfolio has to have the same order as the optimal joint TE and weight portfolio, $w_{cp}^{*}$. We have the following theorem that shows it is possible to estimate a portfolio with weight and TE constraints in large dimensions. Also, consistent estimation of Sharpe Ratio is possible for $p>T$.
Theorem 2
@@
(i).
$\|\hat{\boldsymbol{w}}_{R}-\boldsymbol{w}_{R}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*}\|_{1}=O_{p}(r_{w1})=O_{p}(\bar{s}l_{T}K^{3/2})=o_{p}(1).$
@@
$\left|\frac{\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
@@
$\left|\left(\frac{\widehat{SR}_{R}}{SR_{R}}\right)^{2}-1\right|=O_{p}(r_{w1}K^{5/2})=O_{p}(\bar{s}l_{T}K^{4})=o_{p}(1).$
Remarks. 1. Comparing Theorem 1 with Theorem 2 shows that the rates with joint TE and weight constraints are the same. So upper bounds on estimation errors asymptotically behave the same as in TE and joint TE plus weight constraints. But (9) shows that additional terms are estimated in joint TE plus weight-constrained portfolio compared to TE constrained portfolio; however, asymptotically, the differences disappear. Remark 4 below compares only TE and only weight-constrained portfolios to see the difference.
@@
2. We can also have an estimate for TE as $\widehat{TE}_{cp}:=\sqrt{\hat{\boldsymbol{w}}_{cp}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{cp}}$, as an out-of-sample estimate. It can be shown that $|(\frac{\widehat{TE}_{cp}}{TE})^{2}-1|$ will converge in probability to zero by following the proof of Theorem 2(ii), and the rate will be the same as in Theorem 2(ii). This estimator is consistent.
@@
3. There is also an issue of how TE constraints, jointly with weight constraints, affect the returns. Information Ratio (IR) is used to measure the effect of TE constraints on expected returns; however, Bajeux-Besnainou et al. (2011) shows that IR in the case of weights joint with TE constraints is not an appropriate measure since it can decrease artificially with an increase in TE. They propose Adjusted Information Ratio (AIR). In (22) of Bajeux-Besnainou et al. (2011), it is clear that AIR is measuring the returns of optimal portfolio minus constrained minimum TE portfolio satisfying the weight constraints with respect to risk of that portfolio. We will not be pursuing such an approach since our interest centers on optimal portfolios as in Theorem 2.
4. The case of only weight constraints is analyzed after the proof of Theorem 2 in Appendix A.6 in detail, and the rates are the same as in Theorem 2. Comparing TE only restrictions with only weight restriction portfolio is done in detail in Appendix A.6.1. To summarize, since TE has a benchmark portfolio and it does not have to be estimated, it will incur less estimation errors in $l_{1}$ norm for weights compared with only weight restricted portfolio in finite samples, but asymptotically upper bound estimation error rates are the same as $r_{w1}$.
5. Comparing unconstrained portfolio with joint TE and weight-constrained portfolio is shown via proofs in Appendix A.7 and A.7.3. In finite samples, it is not clear whether joint TE plus weight or unconstrained portfolio estimation is easier in $l_{1}$ norm estimation error-sense. However, asymptotically, upper bounds in estimation errors in both portfolios are the same and $r_{w1}$. Variance and Sharpe Ratio estimation errors behave like weight estimation errors.
6. Remark 7 in Theorem 1 applies here as well. We use diverging $\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}$ in our proofs, but if we impose bounded $\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}$, the consistency is only achieved when $p<T$. This can be seen in our proofs for the bounded Sharpe Ratio case in Appendix B, as in Theorem B.2, Theorem B.4 there.
7. Note that we use same $\kappa$, the risk tolerance parameter both in this section with joint TE and weight constraints and only TE constraint in the previous section. The reason is that $\kappa$ is tied to the TE constraint in both cases, and hence use the same Lagrange multiplier in the constrained problem in p.299, and p.304 of Bajeux-Besnainou et al. (2011). However, the cases of unconstrained portfolio and only weight constrained portfolio is different. There our risk tolerance parameter is $\kappa_{w}$, defined in Appendix A.6 and after (A.116). In cases of unconstrained portfolio and only weight constrained portfolio, $\kappa_{w}$ is tied to the variance constraint in the portfolio, hence $\kappa\neq\kappa_{w}$. But we argue that before Lemma A.3, this does not change the proofs so we assume $\kappa=\kappa_{w}$ to simplify the notation in our proofs. Consistent estimation of $\kappa_{w}$ is possible and can be shown as in Section B.3 for $\kappa$.
5 TE with Inequality Constraints on the Weights of the Portfolio
@@
In practice, inequality constraints on the weight of the portfolio are more common. Of course, one major issue is that how the inequality weight interacts with large number of assets in the portfolio? In this part, we assume $\omega\geq 0$; by that restriction, we exclude the case where the benchmark portfolio, $\boldsymbol{m}$, does not satisfy the inequality constraint. So we exclude an unlikely event in practice. The optimization problem is:
@@
Note that when the weight constraint is not binding, we should get TE solution $\boldsymbol{w}_{d}^{*}$ in Section 3. Otherwise, we should get the solution in Section 4, which is $\boldsymbol{w}_{cp}^{*}$. This $\boldsymbol{w}_{cp}^{*}$ is the joint TE and equality weight constraint based portfolio. We analyze the cases that may lead to binding or nonbinding weight constraint solutions now. We need $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}^{*}\geq\omega$ to have a binding solution. Equivalently, the weight constraint is binding when $\kappa w_{u}\geq\omega$ since $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}_{d}^{*}=\kappa w_{u}$ by (3) and (8). [^1] As in Bajeux-Besnainou et al. (2011) we use greater than equal constraint for binding rather than just equality in Section 4. The nonbinding case will be $\kappa w_{u}<\omega$, as also seen in (17) of Bajeux-Besnainou et al. (2011).
@@
We provide the following theoretical portfolio $\boldsymbol{w}_{op}^{*}:=\boldsymbol{w}_{d}^{*}1_{\{\kappa w_{u}<\omega\}}+\boldsymbol{w}_{cp}^{*}1_{\{\kappa w_{u}>\omega\}}.$ $1_{\{\cdot\}}$ is an indicator function. This means, if the weight constraint is not binding, i.e. $\kappa w_{u}<\omega$, then the portfolio is $\boldsymbol{w}_{d}^{*}$, which is TE (only) constrained portfolio. If the weight constraint is binding, i.e. $\kappa w_{u}>\omega$, then the portfolio is given by joint TE and weight constraint portfolio, which is $\boldsymbol{w}_{cp}^{*}$. For technical reasons, we exclude $\kappa w_{u}=\omega$. The feasible estimator of this theoretical portfolio is given by $\hat{\boldsymbol{w}}_{op}:=\hat{\boldsymbol{w}}_{d}1_{\{\kappa\hat{w}_{u}<\omega\}}+\hat{\boldsymbol{w}}_{cp}1_{\{\kappa\hat{\boldsymbol{w}}_{u}>\omega\}}.$
@@
By (4) and (10), the estimated portfolio with benchmarks taken into account is defined as $\hat{\boldsymbol{w}}_{est}:=\hat{\boldsymbol{w}}1_{\{\kappa\hat{\boldsymbol{w}}_{u}<\omega\}}+\hat{\boldsymbol{w}}_{R}1_{\{\kappa\hat{\boldsymbol{w}}_{u}>\omega\}}.$ The theoretical counterpart is given by  [^2] An equivalent way of writing the nonbinding constraint in the indicator function is by $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}<\omega+\boldsymbol{1}_{R}^{\prime}\boldsymbol{m}$, and the binding constraint in the indicator function can be written as $\boldsymbol{1}_{R}^{\prime}\boldsymbol{w}>\omega+\boldsymbol{1}_{R}^{\prime}\boldsymbol{m}$. $\boldsymbol{w}_{est}^{*}:=\boldsymbol{w}^{*}1_{\{\kappa w_{u}<\omega\}}+\boldsymbol{w}_{R}^{*}1_{\{\kappa w_{u}>\omega\}}.$ We define the Sharpe Ratio whether we are in the binding or nonbinding weights constraint case as follows $SR_{est}^{*}:=\frac{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu}}{\sqrt{\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}}}1_{\{\kappa w_{u}<\omega\}}+\frac{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu}}{\sqrt{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}}1_{\{\kappa w_{u}>\omega\}},$ and its estimator as $\widehat{SR}_{est}:=\frac{\hat{\boldsymbol{w}}^{\prime}\boldsymbol{\mu}}{\sqrt{\hat{\boldsymbol{w}}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}}}1_{\{\kappa\hat{w}_{u}<\omega\}}+\frac{\hat{\boldsymbol{w}}_{R}^{{}^{\prime}}\boldsymbol{\mu}}{\sqrt{\hat{\boldsymbol{w}}_{R}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}}1_{\{\kappa\hat{w}_{u}>\omega\}}.$
@@
Next theorem shows that our weight, variance, and Sharpe Ratio estimates are consistent, regardless of whether inequality constraints are binding or not. This is a new result in the literature. In reality, we do not know whether the theoretical constraints are binding or not. However, our data-based method selects whether the constraints are binding or not and our estimates converge in probability to the optimal weight, variance or Sharpe Ratio. Let $C,c$ be some positive constants.
Theorem 3
@@
(i). If $\kappa w_{u}<\omega$, (non-binding weight restriction) we have the same rate of convergence
$\|\hat{\boldsymbol{w}}_{est}-\boldsymbol{w}_{est}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{op}-\boldsymbol{w}_{op}^{*}\|_{1}=\|\hat{\boldsymbol{w}}-\boldsymbol{w}^{*}\|_{1}=O_{p}(r_{w1})=O_{p}(\bar{s}l_{T}K^{3/2})=o_{p}(1),$
otherwise if $\kappa w_{u}>\omega$, (binding weight restriction) then
$\|\hat{\boldsymbol{w}}_{est}-\boldsymbol{w}_{est}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{op}-\boldsymbol{w}_{op}^{*}\|_{1}=\|\hat{\boldsymbol{w}}_{R}-\boldsymbol{w}_{R}^{*}\|_{1}=O_{p}(r_{w1})=O_{p}(\bar{s}l_{T}K^{3/2})=o_{p}(1).$